### v1

# Fuzzy Information Evolution with Three-Way Decision in Social Network Group Decision-Making

Abstract
In group decision-making (GDM) scenarios, uncertainty, dynamic social structures, and vague information present major challenges for traditional opinion dynamics models. To address these issues, this study proposes a novel social network group decision-making (SNGDM) framework that integrates three-way decision (3WD) theory, dynamic network reconstruction, and linguistic opinion representation. First, the 3WD mechanism is introduced to explicitly model hesitation and ambiguity in agent judgments, thereby preventing irrational decisions. Second, a connection adjustment rule based on opinion similarity is developed, enabling agents to adaptively update their communication links and better reflect the evolving nature of social relationships. Third, linguistic terms are used to describe agent opinions, allowing the model to handle subjective, vague, or incomplete information more effectively. Finally, an integrated multi-agent decision-making framework is constructed, which simultaneously considers individual uncertainty, opinion evolution, and network dynamics. The proposed model is applied to a multi-UAV cooperative decision-making scenario, where simulation results and consensus analysis demonstrate its effectiveness. Experimental comparisons further verify the advantages of the algorithm in enhancing system stability and representing realistic decision-making behaviors.
Index Terms:
Social network group decision-making (SNGDM), Three-way decision (3WD), Consensus analysis.
I Introduction
In multi-agent decision-making, information is often highly uncertain. Traditional centralized information fusion and decision-making systems often face problems such as isolated information and poor coordination. To improve local assessment capabilities and overall situation awareness, distributed information fusion and decision-making mechanisms based on multi-agent systems have attracted increasing attention in recent years [1].
Opinion dynamics models have been widely used as a fundamental framework for distributed information fusion in multi-agent systems. By simulating how individual agents update their information, these models effectively show collective behavior and the process of consensus formation in uncertain environment [2]. In [3], the DeGroot and Hegselmann–Krause (HK) models were combined to propose a hybrid opinion dynamics (HOD) model in the large-scale group decision-making (LSGDM) framework. In [4], a self-confidence evolution model, which encompassed the self-confidence levels of one’s group mates and the passage of time, was proposed. To consider the trust relationship in social network group decision making (SNGDM), a new discount-based uninorm trust propagation operator with 2-tuple linguistic was presented in [5]. Following the idea of the personalized individual semantics (PIS) model, the PIS-based linguistic opinions dynamics (PIS-LOD) model was proposed within the framework of bounded confidence effects [6]. In [7], a novel opinion dynamics model called interval opinion dynamics with dynamic bounded confidence was proposed, in which agents expressed their opinions in numerical intervals and the bounded confidences varied within a specified range over time. By taking different numerical interval opinions and uncertainty tolerances into account, a numerical interval opinion dynamics model was proposed to investigate the process of forming collective opinions in a group of interacting agents [8]. Inspired by the DeGroot model, a novel group decision-making (GDM) model with opinions evolution was established, and a consensus model with minimum adjustments was provided to obtain the optimal adjusted initial opinions and collective consensus opinion [9]. Besides, an opinion similarity mixed (OSM) model and a structural similarity mixed (SSM) model were proposed in [10], in which the strong and weak relations between individuals were considered. Based on the social network DeGroot (SNDG) model, the authors proved that the speed of consensus reaching was subject to the largest self-confidence level of opinion followers and the speed of consensus reaching was subject to the top two self-confidence levels of opinion leaders [11]. In [12], a conflict resolution method considering the trust propagation, conflict detection, and alternatives selection was proposed.
As shown in the above analysis, opinion dynamics models have been widely applied across various fields. However, most of these models assume that agents must always make definite choices. For example, in the HK model, if the opinion difference between agents exceeds a given threshold, they will not interact or update their information. However, in many cases, the actual opinion difference may be only slightly larger than the threshold, yet it completely blocks information exchange, which is often unrealistic in practical situations. To overcome this limitation, this study incorporates the three-way decision (3WD) theory. 3WD has demonstrated unique advantages in various decision-making environments. In [13], the update problem of 3WD with dynamic variation of scales in incomplete multi-scale information systems was studied, where the updating mechanisms of decision granules induced by the similarity relation were exploited through cut refinement and coarsening via attribute value taxonomies. Considering that the existing 3WD models could not effectively handle incomplete multiattribute decision-making (MADM) problems in real life, the preference of decision-makers for each alternative was considered and the concept of predecisions was introduced, whereby an incomplete fuzzy decision system (IFDS) was obtained in [14]. In [15], a kind of 3WD-MADM with probabilistic dominance relations was proposed, where the two state sets were developed by virtue of probabilistic dominance classes. Also, a probabilistic linguistic 3WD method based on regret theory (RT), namely PL-TWDR, was proposed for MADM problems with probabilistic linguistic term sets (PLTSs), considering the effective rationality of a decision-maker in complex decision environments [16]. In [17], the concepts of positive, neutral, and negative alliances with two thresholds were first provided, and then three-way conflict analysis based on the Bayesian minimum risk theory was conducted, where examples were explored to show how the positive, neutral, and negative alliances were computed with a Pythagorean fuzzy loss function given by an expert.
Moreover, the structure of social network is also essential in GDM. In [18], a cooperation index was formulated to identify the non-cooperative behavior, based on which a non-cooperative behavior transformation method was investigated and a bidirectional feedback mechanism was provided for SNGDM. In [19], trust risks in the conflict-eliminating process (CEP) of SNGDM were addressed through third-party monitoring, and a trust risk analysis-based conflict-eliminating model for SNGDM was developed. In [20], an interaction indicator was exploited to represent peer interaction effort (PIE), individual social cooperation networks (ISCNs) were constructed with the log-sigmoid transition technique, and a novel game-theoretic expert importance evaluation model guided by cooperation effects was finally proposed. To address the challenge that decision-makers in a community setting usually exhibited complex social preferences and intricate social interactions, a minimum cost consensus-based SNGDM approach considering altruism-fairness preferences and ordered trust propagation was designed in [21].
Despite the wide application of opinion dynamics models in distributed decision-making tasks, several key limitations remain. First, many classical models such as the DeGroot and HK frameworks assume that agents must always make definite decisions, typically accepting or rejecting a certain view. This binary assumption overlooks the hesitation and ambiguity that often arise in real-world environments, especially when information is incomplete or conflicting. Although the HK model introduces a confidence threshold to support selective interactions, the results are highly sensitive to small changes in this parameter, which reduces its robustness in dynamic settings. Second, most existing models do not incorporate a mechanism to explicitly handle indecision or uncertainty during the opinion updating process. As a result, agents are often forced to take positions even when the available information does not justify a clear choice. This can lead to premature convergence or extreme polarization. Third, the role of network structure is often simplified or treated as static, whereas in many practical scenarios, communication and influence networks evolve dynamically. Agents do not interact with all others equally; rather, their connections change based on factors such as opinion similarity, trust, or communication cost. Neglecting these structural dynamics limits the model’s ability to reflect realistic social interactions and adaptive coordination.
Motivated by the challenges identified above, this paper proposes a novel information updating model. The main contributions of this study are as follows:
  1. 1.
    Introducing 3WD to describe uncertainty in agent judgments. Unlike traditional opinion dynamics models that use binary decisions, this study applies the 3WD mechanism (“accept,” “reject,” and “defer”) to better handle ambiguous or conflicting information. This helps prevent irrational decisions and premature consensus.
  2. 2.
    Combining opinion dynamics with a dynamic social network reconstruction mechanism. A connection adjustment rule based on opinion similarity is proposed, allowing agents to adaptively update their communication links. This reflects the dynamic and selective nature of real-world social interactions.
  3. 3.
    Using linguistic information to better match real-world scenarios. Instead of using only numerical values, agent opinions are described with linguistic terms. This improves the model’s ability to deal with vague, subjective, or incomplete information, making the decision-making process more practical to explain.
  4. 4.
    Developing an integrated decision-making framework for multi-agent coordination. The framework considers uncertainty in individual decisions, opinion evolution in groups, and changes in network structure, which improves the system’s performance in complex environments.
The paper’s structure is as follows: In Section II, the relevant concepts are introduced. In Section III, the new model is proposed. In Section IV, a series of examples are presented. In Section V, the comparison is conducted. The conclusion is given in Section VI.
II Preliminaries
II-A Linguistic Term Set
Definition 1: Let $H=\{h_{i}|i=0,1,\cdots,2\Phi,|\Phi\in N^{*}\}$ be finite and completely ordered linguistic term sets (LTSs). $h_{i}$ represents a possible value for a linguistic variable. $h_{i}$ and $h_{j}$ should satisfy [22, 23]:
(1) $h_{i}\leq h_{j}$ if and only if $i\leq j$;
(2) The negation operation $neg(h_{i})=h_{j}$ if $i+j=2\Phi$;
(3) If $i\geq j$, then $\max\{h_{i},h_{j}\}=h_{i}$;
(4) If $i\geq j$, then $\min\{h_{i},h_{j}\}=h_{j}$.
II-B Linguistic Scale Functions (LSFs)
[…]
$\theta_{j}=\left\{\begin{array}[]{rcl}\frac{a^{t}-a^{t-j}}{2a^{t}-2},&&{0\leq j\leq t}\\ \frac{a^{t}+a^{j-t}-2}{2a^{t}-2},&&{t<j\leq 2t}\end{array}\right.$ (1)
II-C DeGroot Model
Definition 3 [25]: In the DeGroot model, the agents unconditionally trust the other agents. Suppose that $E=\{e_{1},e_{2},\cdots,e_{N}\}$ indicates the agents. $t\in\{1,2,\cdots,T\}$ indicates the discrete variable representing the number of rounds of opinion evolution. For $i\in N$, $y_{i}(t)\in[0,1]$ is the numeric opinion at time $t$. The weight offered by the agent $e_{i}$ to $e_{j}$ is expressed as $\omega_{j}$, where $0\leq\omega_{j}\leq 1$ and $\sum_{j=1}^{N}\omega_{j}=1$ for all $i\in N$. The opinions evolution rule is:
$y_{i}(t+1)=\sum_{j=1}^{N}\omega_{j}y_{j}(t).$ (2)
II-D HK Bounded Confidence Model
Definition 4 [26, 27]: Suppose that $E=\{e_{1},e_{2},\cdots,e_{k}\}$ indicates the agents. For $e_{i}\in E$, $x_{i}(t)\in[0,1]$ is the numeric opinion at time $t$. $X=\{x_{1}(t),x_{2}(t),\cdots,x_{k}(t)\}$ is the opinions of agents express at time $t$. $\varepsilon$ indicates the confidence threshold value. If all agents hold the same confidence threshold $\varepsilon$, the HK model is homogeneous; otherwise, it is heterogeneous.
Step 1: Determine the confidence set.
Let $I(e_{i},X(t))$ be the confidence set of $e_{i}$ at time $t$.
$I(e_{i},X(t))=\{e_{j}||x_{i}(t)-x_{j}(t)|\leq\varepsilon,e_{j}\in E\},\;\;i=1,2,\cdots,k$ (3)
Step 2: Calculate the weights.
When the agent $e_{i}$ update the opinion at time $t$, the weight of the agent $e_{j}$ is:
[…]
where #$\#I(e_{i},Y(t))$ is the cardinality of $I(e_{i},X(t))$.
Step 3: Opinion evolution.
Let $x_{i}(t+1)$ indicates the opinion of $e_{i}$ at time $t+1$, then
$x_{i}(t+1)=w_{i1}x_{1}(t)+w_{i2}x_{2}(t)+\cdots+w_{ik}x_{k}(t)$ (5)
II-E Social Network
Definition 5 [28, 29, 30]: A social network is $G(V,E)$. $V=\{v_{t},\;t\in M\}$ expresses vertices. $E=\{e_{ts},\;t,s\in M\}$ indicates the edges. The edge $e_{ts}=(v_{t},v_{s})$ indicates the relationship between $v_{t}$ and $v_{s}$. $F=(f_{ts})_{m\times m}$ can describe the social network.
$f_{ts}=\left\{\begin{array}[]{rcl}1,&&{(v_{t},v_{s})\in E}\\ 0,&&{otherwise}\end{array}\right.$ (6)
Definition 6 [31, 32]: For $G(V,E)$, the following concepts are introduced.
Density of social network:
It is the ratio of the number of direct pairwise relationships to the possible number of direct relationships.
In-degree $C_{id}(v_{t})$ of the vertex $v_{t}$:
It is employed to describe the reputation of a node, which reflects the integration force.
Out-degree $C_{od}(v_{t})$ of the vertex $v_{t}$:
It is used to describe the communicative nature of a node, which reflects the radiation.
The final centrality degree is $\gamma=\frac{C_{id}(v_{t})+C_{od}(v_{t})}{2(m-1)}$.
II-F 3WD Model
Definition 7 [33]: The 3WD model includes two states $X=\{C,\neg C\}$ and three actions $H=\{h_{A},h_{D},h_{R}\}$, where $C$ and $\neg C$ represent for a “good” state and a “bad” state, respectively. $h_{A}$, $h_{D}$, and $h_{R}$ represent the three domains in classifying an alternative $a_{i}$, that is, $a_{i}\in POS(C),\,a_{i}\in BOU(C),\,\text{and}\;a_{i}\in NEG(C)$, respectively. Furthermore, $POS(C),\,BOU(C),\,\text{and}\;NEG(C)$ correspond to the positive domain, the boundary domain and the negative domain, respectively. When $a_{i}\in C$, let $\lambda_{AP}$, $\lambda_{DP}$, and $\lambda_{RP}$ denote the losses of taking the actions $h_{A}$, $h_{D}$, and $h_{R}$. Similarly, when $a_{i}\in\neg C$, let $\lambda_{AN}$, $\lambda_{DN}$, and $\lambda_{RN}$ represent the losses of taking the actions $h_{A}$, $h_{D}$, and $h_{R}$. Then, the loss functions of $a_{i}$ are given in Table I, where $C(P)$ and $\neg C(P)$ are represented by $P$ and $N$ standing for $C$ and $\neg C$, respectively. For any alternative $a_{i}\in A$, the equivalence class of $a_{i}$ is represented as $[a_{i}]$. $Pr(C\mid[a_{i}])$ and $Pr(\neg C\mid[a_{i}])$ represent the conditional probability of the alternative $a_{i}$ belongs to and not belonging to $C$, respectively. Then, the expected losses $R(h_{I}\mid[a_{i}])$ for the three actions ($I=A;D;R$) are:
$R(h_{A}\mid[a_{i}])=\lambda_{AP}Pr(C\mid[a_{i}])+\lambda_{AN}Pr(\neg C\mid[a_{i}]);$
$R(h_{D}\mid[a_{i}])=\lambda_{DP}Pr(C\mid[a_{i}])+\lambda_{DN}Pr(\neg C\mid[a_{i}]);$
$R(h_{R}\mid[a_{i}])=\lambda_{RP}Pr(C\mid[a_{i}])+\lambda_{RN}Pr(\neg C\mid[a_{i}]).$
TABLE I: Loss Functions of $a_{i}$.
       $C(P)$        $\neg C(N)$
       $h_{A}$        $\lambda_{AP}$        $\lambda_{AN}$
       $h_{D}$        $\lambda_{DP}$        $\lambda_{DN}$
       $h_{R}$        $\lambda_{RP}$        $\lambda_{RN}$
Afterwards, we can obtain the following rules by using the Bayesian theory of minimum losses.
$(P)\;\text{If}\,R(h_{A}\mid[a_{i}])\leq R(h_{D}\mid[a_{i}])\,\text{and}\;R(h_{A}\mid[a_{i}])\leq R(h_{R}\mid[a_{i}]),$
$\text{decide}\;a_{i}\in POS(C);$
$(B)\;\text{If}\,R(h_{D}\mid[a_{i}])\leq R(h_{A}\mid[a_{i}])\,\text{and}\;R(h_{D}\mid[a_{i}])\leq R(h_{R}\mid[a_{i}]),$
$\text{decide}\;a_{i}\in BOU(C);$
$(N)\;\text{If}\,R(h_{R}\mid[a_{i}])\leq R(h_{A}\mid[a_{i}])\,\text{and}\;R(h_{R}\mid[a_{i}])\leq R(h_{D}\mid[a_{i}]),$
$\text{decide}\;a_{i}\in NEG(C).$
III Linguistic 3WD Opinion Dynamics Model over
Dynamic Social Networks
To better model the interaction between opinion evolution and network structure in GDM, this paper proposes a co-evolutionary model that combines the 3WD mechanism with a dynamic network update process. Traditional models often use a binary decision rule, where agents either fully accept or completely reject neighbors based on a fixed confidence bound. However, such a hard-boundary approach cannot fully reflect the diversity of real-world social behavior. In this work, we introduce the 3WD mechanism, which divides the opinion distance between two agents into three regions: if the distance is less than a strict acceptance threshold $\alpha$, the neighbor is deterministically accepted; if the distance exceeds a rejection threshold $\beta$, the neighbor is rejected; and if the distance falls between $\alpha$ and $\beta$, the agent accepts the neighbor with a certain probability. This probability is controlled by an exponential decay function $p=\exp(-\lambda(d_{ij}-\alpha))$, where $d_{ij}$ is the opinion distance and $\lambda$ is a decay factor. This design reflects the psychological tendency that people are more likely to accept opinions that are closer to their own, while opinions farther away are less likely to be accepted. The 3WD mechanism allows the model to capture hesitation, selective trust, and partial agreement, which are common in real social interactions. In addition, to break the limitation of a static network, we introduce a dynamic update mechanism based on cognitive similarity. This mechanism allows the connections between agents to change over time, depending on their current opinions. Specifically, if two agents are not connected but their opinion distance is smaller than a threshold $\delta_{\text{add}}$, a link is added between them with probability $p_{\text{add}}$. On the other hand, if two connected agents show a large opinion difference that exceeds $\delta_{\text{cut}}$, their link may be removed with probability $p_{\text{cut}}$. In this way, the network structure evolves according to the cognitive relationships among agents.
We consider a group of $N$ agents whose opinions are expressed using a finite linguistic term set $H=\{h_{0},h_{1},\dots,h_{2\Phi}\}$. Each term $h_{j}\in H$ is mapped to a real-valued scalar in the interval $[0,1]$ via a predefined nonlinear function $f:H\rightarrow[0,1]$. Let $x_{i}(t)\in H$ denote the linguistic opinion of agent $i$ at time $t$, and let $\theta_{i}(t)=f(x_{i}(t))$ be the corresponding numerical representation.
Notation:
  • $H=\{h_{0},h_{1},\dots,h_{2\Phi}\}$: Linguistic term set
  • $f(h_{j})=\theta_{j}$: Mapping from $h_{j}$ to $[0,1]$
  • $x_{i}(t)\in H$: Linguistic opinion of agent $i$ at time $t$
  • $\theta_{i}(t)=f(x_{i}(t))$: Numerical opinion
  • $A_{ij}(t)$: Social adjacency matrix ($1$ if connected, 00 otherwise)
  • $\alpha$, $\beta$: Three-way decision thresholds
  • $\lambda$: Probability decay parameter
  • $\delta_{\text{add}},\delta_{\text{cut}}$: Network update distance thresholds
  • $p_{\text{add}},p_{\text{cut}}$: Rewiring probabilities
Initialization
Each agent is randomly assigned an initial linguistic opinion $x_{i}(0)\in H$, which is then mapped to its corresponding numeric value $\theta_{i}(0)=f(x_{i}(0))$. The initial social network is represented by an adjacency matrix $A(0)$, where $A_{ij}(0)=1$ if agents $i$ and $j$ are connected, and $A_{ij}(0)=0$ otherwise.
Input:
Linguistic term set $H$, mapping function $f$, initial network $A(0)$, parameters $\alpha,\beta,\lambda,\delta_{\text{add}},\delta_{\text{cut}},p_{\text{add}},p_{\text{cut}},T_{\max},\varepsilon$
Output: Final linguistic opinions $x_{i}$ for all agents
1
2Initialization:
foreach agent $i\in\{1,\dots,N\}$ do
3       Randomly assign $x_{i}(0)\in H$;
4       Compute $\theta_{i}(0)=f(x_{i}(0))$;
5
6
7for $t=0$ to $T_{\max}$ do
8       foreach agent $i\in\{1,\dots,N\}$ do
9             Initialize neighbor set $\mathcal{N}_{i}(t)\leftarrow\emptyset$;
10             foreach $j\neq i$ and $A_{ij}(t)=1$ do
11                   Compute $d_{ij}=|\theta_{i}(t)-\theta_{j}(t)|$;
12                   if $d_{ij}\leq\alpha$ then
                         $\mathcal{N}_{i}(t)\leftarrow\mathcal{N}_{i}(t)\cup\{j\}$ ;
                          // Accept
13
14                  else
15                        if $d_{ij}\geq\beta$ then
                               continue ;
                                // Reject
16
17                        else
18                               Compute $p=\exp(-\lambda(d_{ij}-\alpha))$;
19                               Sample $r\sim\mathcal{U}(0,1)$;
20                               if $r<p$ then
                                     $\mathcal{N}_{i}(t)\leftarrow\mathcal{N}_{i}(t)\cup\{j\}$ ;
                                      // Accept with uncertainty
21
22
23
24
25            if $\mathcal{N}_{i}(t)\neq\emptyset$ then
26                   $\theta_{i}(t+1)=\frac{1}{|\mathcal{N}_{i}(t)|}\sum\limits_{j\in\mathcal{N}_{i}(t)}\theta_{j}(t)$;
27
28            else
                   $\theta_{i}(t+1)=\theta_{i}(t)$ ;
                    // No update
29
30            Map back to linguistic term:;
31             $x_{i}(t+1)=\arg\min\limits_{h_{j}\in H}|f(h_{j})-\theta_{i}(t+1)|$;
32
33
      // Step 5: Dynamic Network Rewiring
34       foreach pair $(i,j)$ with $i\neq j$ do
35             Compute $d_{ij}(t)=|\theta_{i}(t)-\theta_{j}(t)|$;
36             if $A_{ij}(t)=0$ and $d_{ij}(t)<\delta_{\text{add}}$ then
37                   Sample $r\sim\mathcal{U}(0,1)$;
38                   if $r<p_{\text{add}}$ then
                         Set $A_{ij}(t+1)=A_{ji}(t+1)=1$ ;
                          // Add link
39
40
41            if $A_{ij}(t)=1$ and $d_{ij}(t)>\delta_{\text{cut}}$ then
42                   Sample $r\sim\mathcal{U}(0,1)$;
43                   if $r<p_{\text{cut}}$ then
                         Set $A_{ij}(t+1)=A_{ji}(t+1)=0$ ;
                          // Remove link
44
45
46
47
48      Check convergence:;
49       if $\max\limits_{i}|\theta_{i}(t+1)-\theta_{i}(t)|<\varepsilon$ then
50             break;
51
52
Algorithm 1 Linguistic 3WD Opinion Dynamics over Dynamic Social Network
Opinion Updating
At each discrete time step $t=1,2,\dots,T_{\max}$, every agent $i$ updates its opinion based on the opinions of its neighbors. Specifically, agent $i$ evaluates each connected agent $j$ (i.e., $A_{ij}(t)=1$) by computing the opinion distance $d_{ij}(t)=|\theta_{i}(t)-\theta_{j}(t)|$. Then, a 3WD strategy is applied:
  • If $d_{ij}(t)\leq\alpha$, agent $j$ is accepted as a trusted neighbor.
  • If $d_{ij}(t)\geq\beta$, agent $j$ is rejected.
  • If $\alpha<d_{ij}(t)<\beta$, agent $j$ is accepted with probability $p=\exp(-\lambda(d_{ij}(t)-\alpha))$.
Let $\mathcal{N}_{i}(t)$ denote the set of accepted neighbors of agent $i$ at time $t$. If $\mathcal{N}_{i}(t)$ is nonempty, agent $i$ updates its numerical opinion by averaging:
$\theta_{i}(t+1)=\frac{1}{|\mathcal{N}_{i}(t)|}\sum_{j\in\mathcal{N}_{i}(t)}\theta_{j}(t).$
If $\mathcal{N}_{i}(t)$ is empty, then $\theta_{i}(t+1)=\theta_{i}(t)$ and the opinion remains unchanged. After the numerical update, the agent’s opinion is mapped back to the nearest linguistic term:
$x_{i}(t+1)=\arg\min_{h_{j}\in H}|f(h_{j})-\theta_{i}(t+1)|.$
Dynamic Network Rewiring
Following the opinion update, the social network evolves based on the opinion differences among agents. For each pair of agents $(i,j)$:
  • If $A_{ij}(t)=0$ and $|\theta_{i}(t)-\theta_{j}(t)|<\delta_{\text{add}}$, a link is added with probability $p_{\text{add}}$.
  • If $A_{ij}(t)=1$ and $|\theta_{i}(t)-\theta_{j}(t)|>\delta_{\text{cut}}$, the existing link is removed with probability $p_{\text{cut}}$.
The adjacency matrix is then updated to $A(t+1)$ accordingly.
Termination Condition
The iteration continues until the maximum opinion change between two successive steps is smaller than a convergence threshold $\varepsilon$, i.e.,
$\max_{i}|\theta_{i}(t+1)-\theta_{i}(t)|<\varepsilon,$
or until the maximum number of iterations $T_{\max}$ is reached.
Output
The final linguistic opinion $x_{i}$ for each agent is obtained, representing the converged state or the state at termination.
IV Example
In recent years, with the development of multi-UAV cooperative technology, UAV swarm intelligence systems have demonstrated a high level of autonomy and efficient information-sharing capabilities across various fields. It is worth noting that this study is inspired by the research presented in [34]. As shown in Fig. 1, a group of $n$ UAVs is cooperatively circumnavigating a group of $m$ unknown targets on a horizontal plane. Taking post-earthquake rescue in mountainous areas as an example, multiple UAVs are often deployed to search for survivors across large regions. In this setting, the model based on collaborative decision-making and opinion dynamics is essential. Each UAV acts as an intelligent agent, and shares the linguistic information with the system. As these agents communicate, they update the evaluation information of the situation.
Figure 1: Circumnavigation of $m$ Targets by $n$ UAVs (Red Center Represents the Geometric Centroid of $m$ Targets) [34].
To evaluate the proposed linguistic 3WD opinion dynamics model, we conduct simulations with a fixed number of agents and predefined model parameters. Specifically, the number of agents was set to $N=20$. The linguistic term set consisted of $12\Phi+$ 1 elements, with $\Phi=3$, and the non-linear mapping parameter was set to $a=2$. The 3WD is controlled by two thresholds: the strict acceptance threshold $\alpha=0.3$, and the strict rejection threshold $\beta=0.6$. Agents falling in the uncertain region $(\alpha,\beta)$ were accepted with a probability that decays exponentially with the distance from $\alpha$, governed by a decay factor $\lambda=10$. The opinion inertia parameter was set to $\omega=0$, meaning that agents fully adopt the weighted average of their accepted neighbors’ opinions without retaining their previous value. For the dynamic network update process, we adopted an additive threshold $\delta_{\text{add}}=0.15$ and a cut-off threshold $\delta_{\text{cut}}=0.45$. Links were added or removed with probabilities $p_{\text{add}}=0.5$ and $p_{\text{cut}}=0.5$, respectively. The simulation ran for at most $T_{\max}=10$ time steps, or until convergence was reached, defined by a maximum opinion change less than a tolerance of $\varepsilon=10^{-3}$.
The initial linguistic opinions of the 20 agents at time step $t=0$ are given by:
$x_{1}(0)$ $=h_{0},\quad$ $x_{2}(0)$ $=h_{3},\quad$ $x_{3}(0)$ $=h_{6},\quad$ $x_{4}(0)$ $=h_{0},$
$x_{5}(0)$ $=h_{0},\quad$ $x_{6}(0)$ $=h_{0},\quad$ $x_{7}(0)$ $=h_{5},\quad$ $x_{8}(0)$ $=h_{3},$
$x_{9}(0)$ $=h_{3},\quad$ $x_{10}(0)$ $=h_{3},\quad$ $x_{11}(0)$ $=h_{5},\quad$ $x_{12}(0)$ $=h_{1},$
$x_{13}(0)$ $=h_{0},\quad$ $x_{14}(0)$ $=h_{5},\quad$ $x_{15}(0)$ $=h_{3},\quad$ $x_{16}(0)$ $=h_{0},$
$x_{17}(0)$ $=h_{4},\quad$ $x_{18}(0)$ $=h_{3},\quad$ $x_{19}(0)$ $=h_{4},\quad$ $x_{20}(0)$ $=h_{3}.$
Figure 2: Opinions of All Agents Across Iterations.
Figure 3: Final Linguistic Term for Each Agent.
Fig. 2 shows how the opinions of all agents change over time. We can see that several agents adjusted their opinions clearly in the early steps, which means that there is active communication and belief change at the beginning. Most of these changes happened within the first four iterations. After that, the opinions became more stable. By the fifth iteration, most agents had nearly stopped changing. At the seventh iteration, all opinions stayed the same, which means the system reached a stable state and no more updates are needed. The final linguistic term for each agent is shown in Fig. 3. Fig. 4 presents the evolution of the social network and agent opinions from iteration 0 to 7. In each subplot, nodes represent agents and edges represent social ties. Node colors indicate the linguistic term each agent holds at that time step. It can be seen that agents with similar opinions gradually cluster together as the simulation progresses. From iteration 5 onwards, the network evolves into two densely connected subgroups, each associated with an opinion. Fig. 5 shows how the network structure changes over the iterations. The average degree increases quickly from about 1.9 to nearly 12, which means that agents are building more links with each other as iteration goes on. This suggests that agents are becoming more similar in their opinions, so new connections are more likely under the network update rule. Meanwhile, the number of isolated nodes drops from 12 to 0 in the first two steps and stays at zero after that. This means that all agents are connected to the network early on, and the structure remains fully connected during the process. These results show that the network becomes more connected over time.
Figure 4: Temporal Evolution of the Social Network and Linguistic Opinions.
Figure 5: Network Connectivity Trends During Evolution.
V Comparison, Robustness Analysis, and Complexity Analysis
V-A Impact of Broad Acceptance Thresholds on Consensus Formation
Figure 6: Final Linguistic Term for Each Agent.
In 3WD framework where $[\alpha,\beta]$ is the hesitation region, the choice of thresholds directly influences how likely agents are to exchange information. In Fig. 6, $\alpha$ is set to 0.2 and $\beta$ to 0.4, which defines a narrow acceptance range. Under this setting, only a small number of opinion differences fall within the acceptance or hesitation zones, while most are rejected directly. As a result, agents are more likely to avoid opinion exchange, leading to more isolated updates and a wider spread in final linguistic term indices. This produces a more diverse and polarized results. In contrast, Fig. 3 uses $\alpha=0.3$ and $\beta=0.6$, expanding both the acceptance and hesitation regions. More agent pairs are allowed to interact, either certainly or probabilistically, which leads to smoother opinion integration. Therefore, smaller $\alpha$ and $\beta$ reduce communication opportunities and promote diversity, while larger thresholds encourage interaction and support consensus formation.
In Fig. 7, the thresholds are set to $\alpha=0.5$ and $\beta=0.9$, which significantly enlarges both the hesitation and acceptance regions compared to the previous settings. Under this configuration, only opinion differences greater than 0.9 are directly rejected, so most agent pairs have a chance to exchange information. As shown in the figure, the majority of agents reach a high and consistent linguistic term index $h_{3}$, while only a few remain at the lowest level $h_{0}$. This result indicates that a wider acceptance rule allows more interaction among agents, helping them to integrate their opinions. Compared to smaller values of $\alpha$ and $\beta$, this setup is more likely to support global agreement.
Figure 7: Final Linguistic Term for Each Agent.
V-B Impact of Network Scale on Opinion Evolution and Consensus Formation
Figs. 4, 8, and 9 respectively illustrate the evolution process of the system when the number of agents is 20, 30, and 50, under the proposed model with 3WD mechanism and dynamic social network reconstruction rules. As the number of agents increases, the initially sparse social network gradually evolves into several closely connected subgroups. When $N=20$, the connections between agents are relatively concentrated, and two consensus groups are eventually formed. When $N=30$, the trend of network division becomes stronger, and some agents with extreme opinions form isolated small groups disconnected from the main network. In the large-scale system with $N=50$, although the initial connections are very sparse, the main group gradually absorbs most agents through the self-adjusting structure of the network, and eventually forms a unified group with closer connections and more consistent opinions. This phenomenon fully demonstrates the effectiveness and advantages of the proposed algorithm. As the scale of the social network increases, differences between agents grow, and the uncertainty of connections rises. Traditional static network models are likely to face problems such as opinion polarization, information isolation, and failure to reach consensus. In contrast, our model introduces a dynamic linking mechanism based on opinion distance and a “stay undecided” option from the 3WD theory, which flexibly handles agents near the decision boundary and significantly improves the system’s ability to reach consensus under complex conditions. As shown in the figures, this mechanism not only helps to reduce the trend of division caused by sparse networks and opinion differences but also enhances information flow and coordinated evolution within the group.
Figure 8: Temporal Evolution of the Social Network and Linguistic Opinions.
Figure 9: Temporal Evolution of the Social Network and Linguistic Opinions.
V-C Impact of Three-Way Decision Mechanism on Opinion Dynamics
To evaluate the effect of the proposed 3WD mechanism on the evolution of group opinions, we conducted a comparative study involving four scenarios. Fig. 10 shows a comparison of convergence speed in four different scenarios. Each subplot represents a specific setting, where the horizontal axis is the number of iterations and the vertical axis shows the maximum change in opinions per step, denoted as $\delta_{\text{max}}$. This value is used as the measurement of convergence speed in all four experiments, and it is calculated by the following formula:
$\delta_{\text{max}}(t)=\max_{i}|\theta_{i}(t)-\theta_{i}(t-1)|$
Here, $\theta_{i}(t)$ represents the opinion of agent $i$ at $t$. This metric reflects the largest individual change at each step and helps evaluate how fast or smooth the opinion evolution process is. In the experiments, Scenario 1 and Scenario 3 represent the cases with 40 and 60 agents respectively, where 3WD mechanism is applied using parameters $\alpha=0.3$ and $\beta=0.6$. Scenario 2 and Scenario 4 use the same number of agents but do not include the 3WD mechanism, which is represented by $\alpha=\beta=0.6$, meaning no hesitation zone is added. From the figure, we can see that in Scenario 2 and Scenario 4 (without the mechanism), $\delta_{\text{max}}$ drops quickly, and convergence is reached within 4 and 6 iterations respectively. The process is fast and smooth. In contrast, Scenario 1 and Scenario 3 (with the mechanism) show a slower convergence process, with a more gradual decrease in $\delta_{\text{max}}$ and some fluctuations, especially in Scenario 3. Overall, the 3WD mechanism slows down the speed of opinion convergence, which results in a slower but more stable evolution. This effect becomes more visible when the number of agents increases. These results suggest that the 3WD mechanism achieves a balance between convergence speed and system stability.
Figure 10: Convergence Speed Comparison
Under Different 3WD Settings.
V-D Impact of $\delta_{\text{add}}$ on Opinion Dynamics
Fig. 11 shows the evolution of the social network over time steps $t=0,1,2,3,4,5$. The connection threshold $\delta_{\text{add}}$ in Section Example is increased from 0.15 to 0.35. This higher value for adding links allows agents with larger opinion differences to form new connections, making it easier for the network to reconnect. From the figure, we observe that the initial network is fragmented with several isolated nodes. As time increases, agents rapidly form new connections and the network quickly becomes denser. By $t=3$, the system shows signs of global integration, and by $t=5$, all agents are part of a highly connected network with relatively homogeneous opinions. This result suggests that increasing $\delta_{\text{add}}$ facilitates faster structural convergence, reduces isolation, and promotes opinion merging across the entire network.
Figure 11: Temporal Evolution of the Social Network and Linguistic Opinions.
V-E Impact of $\delta_{\text{cut}}$ on Opinion Dynamics
Fig. 12 illustrates the evolution of the social network over time steps $t=0,1,2,\dots,7$. In this example, we decreased the disconnection threshold $\delta_{\text{cut}}$ from 0.45 to 0.2. This makes it easier for existing links to be removed when the opinion distance between agents exceeds the threshold, increasing the likelihood of network fragmentation. As shown in the figure, although the network starts as a connected structure, it quickly begins to break into separate components. By $t=2$, $e_{3}$, $e_{5}$, and $e_{9}$ become isolated from the main group. This pattern persists through later time steps, where the majority of agents form a dense cluster while the isolated ones remain disconnected. These results indicate that a lower $\delta_{\text{cut}}$ can reduce network stability and reinforce persistent social separation.
Figure 12: Temporal Evolution of the Social Network and Linguistic Opinions.
V-F Complexity Analysis
V-F1 Time Complexity
In each iteration, the proposed model checks all possible neighbor pairs $(i,j)$ to compute opinion distances and determine acceptance. In the worst case, each agent needs to compare with $N-1$ others, resulting in a complexity of $\mathcal{O}(N^{2})$ for this step. The dynamic rewiring step also involves checking all agent pairs and may update the network structure, which again takes $\mathcal{O}(N^{2})$ time. The convergence check requires scanning all agents, which takes $\mathcal{O}(N)$ time. Therefore, each iteration has a total time complexity of $\mathcal{O}(N^{2})$, and with a maximum of $T_{\max}$ iterations, the overall time complexity is $\mathcal{O}(T_{\max}\cdot N^{2})$.
V-F2 Space Complexity
The main memory usage comes from storing the network and agent states. The adjacency matrix takes $\mathcal{O}(N^{2})$ space. Each agent’s numerical opinion $\theta_{i}(t)$ and linguistic opinion $x_{i}(t)$ together take $\mathcal{O}(N)$ space. In addition, the neighbor sets may each contain up to $N$ elements, which leads to a worst-case memory usage of $\mathcal{O}(N^{2})$. Therefore, the total space complexity is $\mathcal{O}(N^{2})$.
In summary, this algorithm models the information evolution with a reasonable level of computational complexity. Although the worst-case time and space complexity are both $\mathcal{O}(N^{2})$, real-world social networks are usually sparse and locally connected, so the actual runtime is often much lower than the theoretical upper bound. Moreover, the model has a clear structure and is easy to implement in parallel.
V-G Potential Applications
V-G1 Brain-Computer Interface (BCI) systems
BCI systems often involve non-invasive brain signals such as electroencephalography (EEG) and functional near-infrared spectroscopy (fNIRS). These signals usually have high variability, strong individual differences, and incomplete information. Traditional classification methods require making clear decisions for all states. However, in real-world applications, the user’s cognitive state may be “uncertain” or “hesitant.” Introducing the 3WD theory allows the brain signal recognition system to produce three types of outputs: “accept,” “reject,” and “defer.” This keeps an intermediate state for handling vague information and helps avoid incorrect triggering. In addition, the mechanism of dynamically adjusting neural information channels can also inspire the design of adaptive models for individual brain region communication.
Figure 13: Description of Brain Regions and EEG-Based Paradigms. (a) Four Regions of the Cortex. (b) Four Other Common EEG-Based Paradigms [35].
V-G2 Opinion Fusion in Smart City Governance
In smart city governance, such as public policy making, intelligent traffic management, and urban risk warning, a large amount of public opinion needs to be collected and integrated. When there are differences in views, uncertain perceptions, and vague opinions, traditional opinion fusion models can easily lead to extreme results or distorted information. The proposed model allows individual opinions to be expressed in natural language and handles uncertain areas through deferred decisions. This makes the opinion fusion process more understandable and more tolerant of ambiguity.
V-G3 Multi-Robot Coordination in Disaster Response and Agricultural Monitoring
In scenarios such as post-disaster search and rescue and agricultural monitoring, multi-robot systems need to make collaborative decisions based on environmental sensing, their own abilities, and task requirements. Traditional assignment strategies often require each robot to immediately accept or reject a task. However, in uncertain environments or when task complexity is unclear, this approach may lead to poor task allocation or wasted resources. By introducing the 3WD theory, robots are allowed to delay task decisions, so they can make better choices when more information becomes available. At the same time, by adjusting the communication network dynamically, robots can improve their cooperation based on task similarity or path overlap, which increases the overall efficiency and robustness of the system.
VI Conclusion
This paper proposes a group opinion update model that combines the 3WD mechanism with a dynamic network update process. The introduced 3WD mechanism sets a “hesitation interval” to expand the traditional binary rule of accepting or rejecting neighbors into a three-way probabilistic rule. This allows an agent to accept information from another agent with a certain probability when their opinion difference lies in an uncertain range. In addition, to overcome the limitations of a static network structure, the model includes a dynamic update mechanism based on cognitive similarity. By setting probabilistic rules for adding or removing links, the social relationships between agents can adjust in real time as their opinions evolve. The proposed algorithm is applied to a multi-UAV cooperative decision-making scenario. By simulating the information exchange and dynamic negotiation among UAVs, we demonstrate that the mechanism is effective in practical distributed intelligent systems. The experimental comparisons further verify the advantages of the proposed algorithm.
As AI agents powered by large language models (LLMs) play an increasingly important role in group decision-making, how to combine the diverse solutions generated by different models has become a key issue. Future research can focus on conflict resolution between models, trust evaluation mechanisms, and integration strategies to improve the overall level of collaborative intelligence in the system. Also, we will add reinforcement learning or adaptive weighting based on past interactions to enable agents to evolve their decision strategies over time.

### v2

@@
 Abstract
-In group decision-making (GDM) scenarios, uncertainty, dynamic social structures, and vague information present major challenges for traditional opinion dynamics models.
+In group decision-making (GDM) scenarios, uncertainty, dynamic social structures, and vague information present challenges for traditional opinion dynamics models.
 To address these issues, this study proposes a novel social network group decision-making (SNGDM) framework that integrates three-way decision (3WD) theory, dynamic network reconstruction, and linguistic opinion representation.
@@
 Second, a connection adjustment rule based on opinion similarity is developed, enabling agents to adaptively update their communication links and better reflect the evolving nature of social relationships.
-Third, linguistic terms are used to describe agent opinions, allowing the model to handle subjective, vague, or incomplete information more effectively.
+Third, linguistic terms are used to describe agent opinions, allowing the model to handle vague and incomplete information more effectively.
 Finally, an integrated multi-agent decision-making framework is constructed, which simultaneously considers individual uncertainty, opinion evolution, and network dynamics.
 The proposed model is applied to a multi-UAV cooperative decision-making scenario, where simulation results and consensus analysis demonstrate its effectiveness.
-Experimental comparisons further verify the advantages of the algorithm in enhancing system stability and representing realistic decision-making behaviors.
+Experimental comparisons further verify the algorithm’s advantages in enhancing system stability and representing realistic decision-making behaviors.
 Index Terms:
@@
 Traditional centralized information fusion and decision-making systems often face problems such as isolated information and poor coordination.
-To improve local assessment capabilities and overall situation awareness, distributed information fusion and decision-making mechanisms based on multi-agent systems have attracted increasing attention in recent years [1].
+To enhance local assessment capabilities and overall situational awareness, distributed information fusion and decision-making mechanisms based on multi-agent systems have garnered increasing attention in recent years [1].
 Opinion dynamics models have been widely used as a fundamental framework for distributed information fusion in multi-agent systems.
-By simulating how individual agents update their information, these models effectively show collective behavior and the process of consensus formation in uncertain environment [2].
+By simulating how individual agents update their information, these models effectively show collective behavior and the process of consensus formation in an uncertain environment [2].
 In [3], the DeGroot and Hegselmann–Krause (HK) models were combined to propose a hybrid opinion dynamics (HOD) model in the large-scale group decision-making (LSGDM) framework.
@@
 Following the idea of the personalized individual semantics (PIS) model, the PIS-based linguistic opinions dynamics (PIS-LOD) model was proposed within the framework of bounded confidence effects [6].
-In [7], a novel opinion dynamics model called interval opinion dynamics with dynamic bounded confidence was proposed, in which agents expressed their opinions in numerical intervals and the bounded confidences varied within a specified range over time.
+In [7], a novel opinion dynamics model called interval opinion dynamics with dynamic bounded confidence was proposed, in which agents expressed their opinions in numerical intervals.
 By taking different numerical interval opinions and uncertainty tolerances into account, a numerical interval opinion dynamics model was proposed to investigate the process of forming collective opinions in a group of interacting agents [8].
-Inspired by the DeGroot model, a novel group decision-making (GDM) model with opinions evolution was established, and a consensus model with minimum adjustments was provided to obtain the optimal adjusted initial opinions and collective consensus opinion [9].
+Inspired by the DeGroot model, a novel group decision-making (GDM) model with opinion evolution was established, and a consensus model with minimum adjustments was provided to obtain the optimal adjusted initial opinions and collective consensus opinion [9].
 Besides, an opinion similarity mixed (OSM) model and a structural similarity mixed (SSM) model were proposed in [10], in which the strong and weak relations between individuals were considered.
-Based on the social network DeGroot (SNDG) model, the authors proved that the speed of consensus reaching was subject to the largest self-confidence level of opinion followers and the speed of consensus reaching was subject to the top two self-confidence levels of opinion leaders [11].
-In [12], a conflict resolution method considering the trust propagation, conflict detection, and alternatives selection was proposed.
+Based on the social network DeGroot model, the authors proved that the speed of consensus reaching was subject to the most significant self-confidence level of opinion followers and the speed of consensus reaching was subject to the top two self-confidence levels of opinion leaders [11].
+In [12], a conflict resolution method considering the trust propagation, conflict detection, and alternative selection was proposed.
 As shown in the above analysis, opinion dynamics models have been widely applied across various fields.
@@
 In [13], the update problem of 3WD with dynamic variation of scales in incomplete multi-scale information systems was studied, where the updating mechanisms of decision granules induced by the similarity relation were exploited through cut refinement and coarsening via attribute value taxonomies.
-Considering that the existing 3WD models could not effectively handle incomplete multiattribute decision-making (MADM) problems in real life, the preference of decision-makers for each alternative was considered and the concept of predecisions was introduced, whereby an incomplete fuzzy decision system (IFDS) was obtained in [14].
+Considering that the existing 3WD models could not effectively handle incomplete multiattribute decision-making (MADM) problems in real life, the preference of decision-makers for each alternative was considered, and the concept of predecisions was introduced, whereby an incomplete fuzzy decision system (IFDS) was obtained in [14].
 In [15], a kind of 3WD-MADM with probabilistic dominance relations was proposed, where the two state sets were developed by virtue of probabilistic dominance classes.
-Also, a probabilistic linguistic 3WD method based on regret theory (RT), namely PL-TWDR, was proposed for MADM problems with probabilistic linguistic term sets (PLTSs), considering the effective rationality of a decision-maker in complex decision environments [16].
-In [17], the concepts of positive, neutral, and negative alliances with two thresholds were first provided, and then three-way conflict analysis based on the Bayesian minimum risk theory was conducted, where examples were explored to show how the positive, neutral, and negative alliances were computed with a Pythagorean fuzzy loss function given by an expert.
-Moreover, the structure of social network is also essential in GDM.
-In [18], a cooperation index was formulated to identify the non-cooperative behavior, based on which a non-cooperative behavior transformation method was investigated and a bidirectional feedback mechanism was provided for SNGDM.
+Also, a probabilistic linguistic 3WD method based on regret theory (RT), namely PL-TWDR, was proposed for MADM problems with probabilistic linguistic term sets (PLTSs), considering the practical rationality of a decision-maker in complex decision environments [16].
+In [17], the concepts of positive, neutral, and negative alliances with two thresholds were first provided.
+Then, a three-way conflict analysis based on the Bayesian minimum risk theory was conducted, where examples were explored to show how the positive, neutral, and negative alliances were computed with a Pythagorean fuzzy loss function given by an expert.
+Moreover, the structure of the social network is also essential in GDM.
+In [18], a cooperation index was formulated to identify the non-cooperative behavior, based on which a non-cooperative behavior transformation method was investigated, and a bidirectional feedback mechanism was provided for SNGDM.
 In [19], trust risks in the conflict-eliminating process (CEP) of SNGDM were addressed through third-party monitoring, and a trust risk analysis-based conflict-eliminating model for SNGDM was developed.
 In [20], an interaction indicator was exploited to represent peer interaction effort (PIE), individual social cooperation networks (ISCNs) were constructed with the log-sigmoid transition technique, and a novel game-theoretic expert importance evaluation model guided by cooperation effects was finally proposed.
-To address the challenge that decision-makers in a community setting usually exhibited complex social preferences and intricate social interactions, a minimum cost consensus-based SNGDM approach considering altruism-fairness preferences and ordered trust propagation was designed in [21].
-Despite the wide application of opinion dynamics models in distributed decision-making tasks, several key limitations remain.
-First, many classical models such as the DeGroot and HK frameworks assume that agents must always make definite decisions, typically accepting or rejecting a certain view.
-This binary assumption overlooks the hesitation and ambiguity that often arise in real-world environments, especially when information is incomplete or conflicting.
-Although the HK model introduces a confidence threshold to support selective interactions, the results are highly sensitive to small changes in this parameter, which reduces its robustness in dynamic settings.
-Second, most existing models do not incorporate a mechanism to explicitly handle indecision or uncertainty during the opinion updating process.
-As a result, agents are often forced to take positions even when the available information does not justify a clear choice.
-This can lead to premature convergence or extreme polarization.
-Third, the role of network structure is often simplified or treated as static, whereas in many practical scenarios, communication and influence networks evolve dynamically.
-Agents do not interact with all others equally; rather, their connections change based on factors such as opinion similarity, trust, or communication cost.
-Neglecting these structural dynamics limits the model’s ability to reflect realistic social interactions and adaptive coordination.
+To address the challenge that decision-makers in a community setting typically exhibit complex social preferences and intricate social interactions, a minimum cost consensus-based SNGDM approach was designed, considering altruism-fairness preferences and ordered trust propagation [21].
+Despite the wide use of opinion dynamics models in distributed decision-making, several important challenges remain.
+First, some classical models, such as the DeGroot and HK frameworks, assume that agents must always make a clear decision, usually to accept or reject a particular opinion, and do not provide a mechanism to express hesitation or uncertainty explicitly.
+This forces agents to take a position even when the available information is not enough, which may lead to premature consensus or extreme polarization.
+For example, in a public policy vote, if a person does not have enough information about the pros and cons of a proposal, they may prefer to stay undecided until more evidence is available.
+However, in many classical models, this person is still forced to move toward “support” or “oppose” in each update step.
+The HK model partly addresses this problem by introducing a confidence threshold to limit interactions, but its results are susceptible to small changes in this threshold.
+For example, when the threshold is set to 0.3, if the opinion difference between two agents is 0.29, they will interact; but if the difference is 0.31, they will refuse to interact.
+In reality, these two situations are almost the same, but the model produces completely different results.
+This “boundary effect” renders the model highly sensitive to threshold settings, which reduces its robustness in dynamic environments.
+Second, some studies treat the network structure as fixed or oversimplified.
+In reality, communication and influence networks change over time.
+Agents interact more frequently with some individuals than others, and these connections evolve in response to factors such as opinion similarity, trust, or communication costs.
+For example, on social media platforms, people may increase interaction if their opinions become closer, or reduce contact and even unfollow, if their views diverge significantly.
+Models that maintain a static network cannot capture this adaptive connection change, which limits their ability to describe fundamental social interactions and coordination.
 Motivated by the challenges identified above, this paper proposes a novel information updating model. The main contributions of this study are as follows:
   1. 1.
-    Introducing 3WD to describe uncertainty in agent judgments.
-Unlike traditional opinion dynamics models that use binary decisions, this study applies the 3WD mechanism (“accept,” “reject,” and “defer”) to better handle ambiguous or conflicting information.
-This helps prevent irrational decisions and premature consensus.
+    Three-way decision (3WD) mechanism for hesitation and uncertainty.
+We design an opinion updating rule based on the three-way decision framework, where each agent can choose “accept,” “reject,” or “defer” in response to incoming information.
+This differs from classical models such as DeGroot or HK that force agents into binary decisions.
+The mechanism explicitly preserves hesitation intervals, reducing the risk of premature consensus and improving decision robustness under conflicting or incomplete information.
   2. 2.
-    Combining opinion dynamics with a dynamic social network reconstruction mechanism.
-A connection adjustment rule based on opinion similarity is proposed, allowing agents to adaptively update their communication links.
-This reflects the dynamic and selective nature of real-world social interactions.
+    Dynamic social network reconstruction driven by opinion similarity.
+We introduce a probabilistic link creation and deletion rule that updates the social network structure according to opinion distances between agents.
+This adaptive adjustment enables agents to strengthen ties with similar peers and weaken ties with divergent ones, capturing the evolving nature of real-world communication patterns.
   3. 3.
-    Using linguistic information to better match real-world scenarios.
-Instead of using only numerical values, agent opinions are described with linguistic terms.
-This improves the model’s ability to deal with vague, subjective, or incomplete information, making the decision-making process more practical to explain.
-  4. 4.
-    Developing an integrated decision-making framework for multi-agent coordination.
-The framework considers uncertainty in individual decisions, opinion evolution in groups, and changes in network structure, which improves the system’s performance in complex environments.
+    Linguistic opinion representation for vague and subjective information.
+We represent opinions using linguistic terms mapped from numerical values, instead of using only numerical states.
+This makes the model better suited for situations where information is subjective, vague, or incomplete.
+The approach improves interpretability and aligns simulation outputs more closely with real-world decision-making contexts, as validated in multi-agent test scenarios.
 The paper’s structure is as follows: In Section II, the relevant concepts are introduced.
@@
 II-B Linguistic Scale Functions (LSFs)
-[…]
+Definition 2 [24]: For the linguistic term $h_{j}$ in $H=\{h_{i}|i=0,1,\cdots,2\Phi,|\Phi\in N^{*}\}$, $h_{j}$ represents a possible value for a linguistic variable.
+If $\theta\in[0,\,1]$ is a numerical value, then the function $f:h_{j}\rightarrow\theta_{j}$, which conducts the mapping from $h_{j}$ to $\theta_{j}\;(j=0,\;1,\;\cdots,\;2t)$, is defined as follows:
 $\theta_{j}=\left\{\begin{array}[]{rcl}\frac{a^{t}-a^{t-j}}{2a^{t}-2},&&{0\leq j\leq t}\\ \frac{a^{t}+a^{j-t}-2}{2a^{t}-2},&&{t<j\leq 2t}\end{array}\right$. (1)
 II-C DeGroot Model
-Definition 3 [25]: In the DeGroot model, the agents unconditionally trust the other agents.
+Definition 3 [25]: In the DeGroot model, the agents unconditionally trust each other.
 Suppose that $E=\{e_{1},e_{2},\cdots,e_{N}\}$ indicates the agents.
@@
 The weight offered by the agent $e_{i}$ to $e_{j}$ is expressed as $\omega_{j}$, where $0\leq\omega_{j}\leq 1$ and $\sum_{j=1}^{N}\omega_{j}=1$ for all $i\in N$.
-The opinions evolution rule is:
+The opinion evolution rule is:
 $y_{i}(t+1)=\sum_{j=1}^{N}\omega_{j}y_{j}(t).$ (2)
@@
 Step 2: Calculate the weights.
-When the agent $e_{i}$ update the opinion at time $t$, the weight of the agent $e_{j}$ is:
+When the agent $e_{i}$ updates the opinion at time $t$, the weight of the agent $e_{j}$ is:
 […]
-where #$\#I(e_{i},Y(t))$ is the cardinality of $I(e_{i},X(t))$.
+where $I(e_{i},Y(t))$ is the cardinality of $I(e_{i},X(t))$.
 Step 3: Opinion evolution.
-Let $x_{i}(t+1)$ indicates the opinion of $e_{i}$ at time $t+1$, then
+Let $x_{i}(t+1)$ indicate the opinion of $e_{i}$ at time $t+1$, then
 $x_{i}(t+1)=w_{i1}x_{1}(t)+w_{i2}x_{2}(t)+\cdots+w_{ik}x_{k}(t)$ (5)
@@
 The edge $e_{ts}=(v_{t},v_{s})$ indicates the relationship between $v_{t}$ and $v_{s}$.
-$F=(f_{ts})_{m\times m}$ can describe the social network.
+$F=(f_{ts})_{m\times m}$ describes the social network.
 $f_{ts}=\left\{\begin{array}[]{rcl}1,&&{(v_{t},v_{s})\in E}\\ 0,&&{otherwise}\end{array}\right$. (6)
@@
 $h_{A}$, $h_{D}$, and $h_{R}$ represent the three domains in classifying an alternative $a_{i}$, that is, $a_{i}\in POS(C),\,a_{i}\in BOU(C),\,\text{and}\;a_{i}\in NEG(C)$, respectively.
-Furthermore, $POS(C),\,BOU(C),\,\text{and}\;NEG(C)$ correspond to the positive domain, the boundary domain and the negative domain, respectively.
+Furthermore, $POS(C),\,BOU(C),\,\text{and}\;NEG(C)$ correspond to the positive domain, the boundary domain, and the negative domain, respectively.
 When $a_{i}\in C$, let $\lambda_{AP}$, $\lambda_{DP}$, and $\lambda_{RP}$ denote the losses of taking the actions $h_{A}$, $h_{D}$, and $h_{R}$.
@@
 Then, the loss functions of $a_{i}$ are given in Table I, where $C(P)$ and $\neg C(P)$ are represented by $P$ and $N$ standing for $C$ and $\neg C$, respectively.
-For any alternative $a_{i}\in A$, the equivalence class of $a_{i}$ is represented as $[a_{i}]$.
+For any alternative $a_{i}\in A$, the equivalence class of $a_{i}$ is represented as $a_{i}]$.
 $Pr(C\mid[a_{i}])$ and $Pr(\neg C\mid[a_{i}])$ represent the conditional probability of the alternative $a_{i}$ belongs to and not belonging to $C$, respectively.
@@
 This design reflects the psychological tendency that people are more likely to accept opinions that are closer to their own, while opinions farther away are less likely to be accepted.
-The 3WD mechanism allows the model to capture hesitation, selective trust, and partial agreement, which are common in real social interactions.
+The 3WD mechanism allows the model to capture hesitation, selective trust, and partial agreement, which are common in fundamental social interactions.
 In addition, to break the limitation of a static network, we introduce a dynamic update mechanism based on cognitive similarity.
@@
 Specifically, if two agents are not connected but their opinion distance is smaller than a threshold $\delta_{\text{add}}$, a link is added between them with probability $p_{\text{add}}$.
-On the other hand, if two connected agents show a large opinion difference that exceeds $\delta_{\text{cut}}$, their link may be removed with probability $p_{\text{cut}}$.
+On the other hand, if two connected agents show a significant opinion difference that exceeds $\delta_{\text{cut}}$, their link may be removed with probability $p_{\text{cut}}$.
 In this way, the network structure evolves according to the cognitive relationships among agents.
@@
   • $\theta_{i}(t)=f(x_{i}(t))$: Numerical opinion
-  • $A_{ij}(t)$: Social adjacency matrix ($1$ if connected, 00 otherwise)
+  • $A_{ij}(t)$: Social adjacency matrix ($1$ if connected, 0 otherwise)
   • $\alpha$, $\beta$: Three-way decision thresholds
@@
 1, a group of $n$ UAVs is cooperatively circumnavigating a group of $m$ unknown targets on a horizontal plane.
-Taking post-earthquake rescue in mountainous areas as an example, multiple UAVs are often deployed to search for survivors across large regions.
+In this paper, we take post-earthquake rescue in mountainous areas as an example.
+Multiple UAVs are often deployed to search for survivors across large regions.
 In this setting, the model based on collaborative decision-making and opinion dynamics is essential.
-Each UAV acts as an intelligent agent, and shares the linguistic information with the system.
+Each UAV acts as an intelligent agent and shares the linguistic information with the system.
 As these agents communicate, they update the evaluation information of the situation.
@@
 Specifically, the number of agents was set to $N=20$.
-The linguistic term set consisted of $12\Phi+$ 1 elements, with $\Phi=3$, and the non-linear mapping parameter was set to $a=2$.
+The linguistic term set consisted of $2\Phi+1$ elements, with $\Phi=3$, and the non-linear mapping parameter was set to $a=2$.
 The 3WD is controlled by two thresholds: the strict acceptance threshold $\alpha=0.3$, and the strict rejection threshold $\beta=0.6$.
@@
 2 shows how the opinions of all agents change over time.
-We can see that several agents adjusted their opinions clearly in the early steps, which means that there is active communication and belief change at the beginning.
+We can see that several agents adjusted their opinions clearly in the early stages, which indicates active communication and belief change at the outset.
 Most of these changes happened within the first four iterations.
@@
 5 shows how the network structure changes over the iterations.
-The average degree increases quickly from about 1.9 to nearly 12, which means that agents are building more links with each other as iteration goes on.
+The average degree increases quickly from about 1.9 to nearly 12, which means that agents are building more links with each other as the iteration goes on.
 This suggests that agents are becoming more similar in their opinions, so new connections are more likely under the network update rule.
-Meanwhile, the number of isolated nodes drops from 12 to 0 in the first two steps and stays at zero after that.
+Meanwhile, the number of isolated nodes drops from 1 to 0 in the first two steps and stays at zero after that.
 This means that all agents are connected to the network early on, and the structure remains fully connected during the process.
 These results show that the network becomes more connected over time.
+To quantitatively evaluate the consensus degree of the system, this paper adopts three complementary indicators.
+First, the opinion variance $\mathrm{Var}(t)=\frac{1}{n}\sum_{i=1}^{n}\bigl(x_{i}(t)-\bar{x}(t)\bigr)^{2}$ is used to reflect the overall dispersion of the group, where a smaller variance indicates stronger convergence.
+Second, the opinion range $R(t)=\max_{i}x_{i}(t)-\min_{i}x_{i}(t)$ is applied to measure the extreme difference within the group, and $R(t)=0$ means that all agents have reached complete consensus.
+Finally, we introduce the normalized consensus index $C_{\mathrm{AAD}}(t)=1-\frac{1}{n}\sum_{i=1}^{n}\frac{|x_{i}(t)-\bar{x}(t)|}{d_{\max}}$ based on the average absolute deviation, and $d_{\max}$ is the maximum possible deviation in the opinion space.
+This index is normalized to the interval $[0,1]$, and values closer to 1 indicate a higher level of consensus.
+By combining these three measures, the consensus level of the system can be comprehensively evaluated from the perspectives of overall convergence, extreme differences, and normalized description.
+In this case, we can obtain $\mathrm{Var}(7)=0.0469$, $R(7)=0.5$, and $C_{\mathrm{AAD}}(7)=0.625$.
+These values show that the opinions still have some dispersion (since the variance 0.0469 is not close to 0), and the largest difference between individuals is 0.5, which means there is a clear gap.
+At the same time, the normalized consensus index is 0.625, which indicates that the system has not reached full agreement but rather a state of partial consensus.
 Figure 4: Temporal Evolution of the Social Network and Linguistic Opinions.
@@
 V Comparison, Robustness Analysis, and Complexity Analysis
-V-A Impact of Broad Acceptance Thresholds on Consensus Formation
+V-A Impact of Parameters and Mechanisms on Proposed Model
+V-A1 Impact of Broad Acceptance Thresholds on Consensus Formation
 Figure 6: Final Linguistic Term for Each Agent.
-In 3WD framework where $[\alpha,\beta]$ is the hesitation region, the choice of thresholds directly influences how likely agents are to exchange information.
+In the 3WD framework, where $\alpha,\beta]$ is the hesitation region, the choice of thresholds directly influences how likely agents are to exchange information.
 In Fig.
@@
 As a result, agents are more likely to avoid opinion exchange, leading to more isolated updates and a wider spread in final linguistic term indices.
-This produces a more diverse and polarized results.
+This produces more diverse and polarized results.
 In contrast, Fig.
@@
 As shown in the figure, the majority of agents reach a high and consistent linguistic term index $h_{3}$, while only a few remain at the lowest level $h_{0}$.
-This result indicates that a wider acceptance rule allows more interaction among agents, helping them to integrate their opinions.
+This result suggests that a wider acceptance rule enables more interaction among agents, facilitating their integration of opinions.
 Compared to smaller values of $\alpha$ and $\beta$, this setup is more likely to support global agreement.
 Figure 7: Final Linguistic Term for Each Agent.
-V-B Impact of Network Scale on Opinion Evolution and Consensus Formation
+V-A2 Impact of Network Scale on Opinion Evolution and Consensus Formation
 Figs.
@@
 When $N=20$, the connections between agents are relatively concentrated, and two consensus groups are eventually formed.
-When $N=30$, the trend of network division becomes stronger, and some agents with extreme opinions form isolated small groups disconnected from the main network.
-In the large-scale system with $N=50$, although the initial connections are very sparse, the main group gradually absorbs most agents through the self-adjusting structure of the network, and eventually forms a unified group with closer connections and more consistent opinions.
+When $N=30$, the trend of network division becomes stronger, and some agents with extreme opinions form isolated small groups disconnected from the leading network.
+In the large-scale system with $N=50$, although the initial connections are very sparse, the leading group gradually absorbs most agents through the self-adjusting structure of the network.
+Eventually, it forms a unified group with closer connections and more consistent opinions.
 This phenomenon fully demonstrates the effectiveness and advantages of the proposed algorithm.
@@
 Figure 9: Temporal Evolution of the Social Network and Linguistic Opinions.
-V-C Impact of Three-Way Decision Mechanism on Opinion Dynamics
+V-A3 Impact of Three-Way Decision Mechanism on Opinion Dynamics
 To evaluate the effect of the proposed 3WD mechanism on the evolution of group opinions, we conducted a comparative study involving four scenarios.
@@
 Each subplot represents a specific setting, where the horizontal axis is the number of iterations and the vertical axis shows the maximum change in opinions per step, denoted as $\delta_{\text{max}}$.
-This value is used as the measurement of convergence speed in all four experiments, and it is calculated by the following formula:
+This value is used as the measurement of convergence speed in all four experiments, and the following formula calculates it:
 $\delta_{\text{max}}(t)=\max_{i}|\theta_{i}(t)-\theta_{i}(t-1)|$
@@
 This metric reflects the largest individual change at each step and helps evaluate how fast or smooth the opinion evolution process is.
-In the experiments, Scenario 1 and Scenario 3 represent the cases with 40 and 60 agents respectively, where 3WD mechanism is applied using parameters $\alpha=0.3$ and $\beta=0.6$.
-Scenario 2 and Scenario 4 use the same number of agents but do not include the 3WD mechanism, which is represented by $\alpha=\beta=0.6$, meaning no hesitation zone is added.
-From the figure, we can see that in Scenario 2 and Scenario 4 (without the mechanism), $\delta_{\text{max}}$ drops quickly, and convergence is reached within 4 and 6 iterations respectively.
+In the experiments, Scenarios 1 and 3 represent cases with 40 and 60 agents, respectively, where the 3WD mechanism is applied using parameters $\alpha=0.3$ and $\beta=0.6$.
+Scenarios 2 and 4 use the same number of agents but do not include the 3WD mechanism, which is represented by $\alpha=\beta=0.6$, meaning no hesitation zone is added.
+From the figure, we can see that in Scenarios 2 and 4 (without the mechanism), $\delta_{\text{max}}$ drops quickly, and convergence is reached within 4 and 6 iterations, respectively.
 The process is fast and smooth.
-In contrast, Scenario 1 and Scenario 3 (with the mechanism) show a slower convergence process, with a more gradual decrease in $\delta_{\text{max}}$ and some fluctuations, especially in Scenario 3.
+In contrast, Scenarios 1 and 3 (with the mechanism) show a slower convergence process, with a more gradual decrease in $\delta_{\text{max}}$ and some fluctuations, especially in Scenario 3.
 Overall, the 3WD mechanism slows down the speed of opinion convergence, which results in a slower but more stable evolution.
@@
 Under Different 3WD Settings.
-V-D Impact of $\delta_{\text{add}}$ on Opinion Dynamics
+V-A4 Impact of $\delta_{\text{add}}$ on Opinion Dynamics
 Fig.
@@
 Figure 11: Temporal Evolution of the Social Network and Linguistic Opinions.
-V-E Impact of $\delta_{\text{cut}}$ on Opinion Dynamics
+V-A5 Impact of $\delta_{\text{cut}}$ on Opinion Dynamics
 Fig.
@@
 Figure 12: Temporal Evolution of the Social Network and Linguistic Opinions.
-V-F Complexity Analysis
-V-F1 Time Complexity
+V-B Complexity Analysis
+V-B1 Time Complexity
 In each iteration, the proposed model checks all possible neighbor pairs $(i,j)$ to compute opinion distances and determine acceptance.
-In the worst case, each agent needs to compare with $N-1$ others, resulting in a complexity of $\mathcal{O}(N^{2})$ for this step.
+In the worst case, each agent needs to compare with N1N-1 others, resulting in a complexity of $\mathcal{O}(N^{2})$ for this step.
 The dynamic rewiring step also involves checking all agent pairs and may update the network structure, which again takes $\mathcal{O}(N^{2})$ time.
@@
 Therefore, each iteration has a total time complexity of $\mathcal{O}(N^{2})$, and with a maximum of $T_{\max}$ iterations, the overall time complexity is $\mathcal{O}(T_{\max}\cdot N^{2})$.
-V-F2 Space Complexity
-The main memory usage comes from storing the network and agent states.
+V-B2 Space Complexity
+The primary source of memory usage is storing the network and agent states.
 The adjacency matrix takes $\mathcal{O}(N^{2})$ space.
@@
 Therefore, the total space complexity is $\mathcal{O}(N^{2})$.
+To test how computation time changes with the number of agents, we carried out experiments on the MATLAB 2020a platform, and the results are shown in Fig.
+13.
+Between 10 and 20 agents, the computation time increases clearly, from about 3.5 seconds to nearly 7.8 seconds.
+When the number of agents rises from 20 to 25, the time drops slightly to about 6.7 seconds.
+After that, between 25 and 35 agents, the time increases again and reaches about 8.2 seconds at 35 agents.
+These results indicate that as the number of agents increases, the overall computation time grows, but the fluctuation is not significant and the growth is not strictly linear.
 In summary, this algorithm models the information evolution with a reasonable level of computational complexity.
 Although the worst-case time and space complexity are both $\mathcal{O}(N^{2})$, real-world social networks are usually sparse and locally connected, so the actual runtime is often much lower than the theoretical upper bound.
-Moreover, the model has a clear structure and is easy to implement in parallel.
-V-G Potential Applications
-V-G1 Brain-Computer Interface (BCI) systems
+Moreover, the model has a clear structure and is easily implementable in parallel.
+Figure 13: Impact of Agent Number on Computation Time.
+V-C Comparative Analysis with Mainstream Models
+In response to the reviewer’s suggestion, we conducted additional comparative experiments with several representative opinion dynamics models.
+The aim was to evaluate the proposed model against widely used approaches and to verify whether the observed behaviors are consistent or significantly different.
+All simulations in this section use the same initial opinions as those described in Section Example.
+V-C1 DeGroot Model
+The DeGroot model is a classical linear consensus model in which each agent updates its opinion by taking a weighted average of all agents’ opinions, as introduced in Definition 3.
+Here, two variants of the DeGroot model are considered:
+  • Uniform weights: In this case, each agent assigns the same weight to every other agent, regardless of the opinion distance. This represents the most basic form of the DeGroot update rule.
+  • Distance-based weights: In this variant, the interaction weight between two agents depends on the distance between their opinions.
+The weight matrix is calculated using $\omega_{j}=\frac{e^{-|D_{ij}|}}{\sum_{k=1}^{20}e^{-|D_{ik}|}}$.
+This formulation ensures that agents with more similar opinions exert a stronger influence on each other.
+Under the DeGroot model, we compare two methods: uniform weights and distance-based weights.
+The results are shown in Fig.
+14.
+In the uniform weight case, the weight of each agent is the same, 1/20.
+Since each row of the weight matrix is identical, the system quickly converges after the first iteration to the arithmetic mean of all initial opinions.
+It remains unchanged, showing an averaging effect and one-step convergence.
+In the distance-based weight case, agents with higher opinion similarity have a stronger influence on each other.
+In the early stage, agents with extreme opinion values change faster, while those in the middle change more slowly.
+The system becomes stable after 2–3 iterations, and the final consensus value is the weighted average of all initial opinions, which is slightly different from the simple average in the uniform weight case.
+These results show that the DeGroot model is sensitive to the way the weights are calculated.
+V-C2 Homogeneous HK Bounded Confidence Model
+The homogeneous HK bounded confidence model assumes that all agents share the same confidence bound $\epsilon$, as introduced in Definition 4.
+At each iteration, an agent only considers the opinions of others whose opinion difference is less than $\epsilon$.
+To enable a thorough comparative analysis, we tested six different values of $\epsilon$, namely $0.35$, $0.30$, $0.25$, $0.20$, $0.15$, and $0.10$.
+As $\epsilon$ decreases, the communication threshold between agents becomes smaller, meaning that an agent interacts only with those whose opinions are very similar to its own.
+Fig.
+15 illustrates the opinion evolution of all agents under six different confidence bound values in the homogeneous HK bounded confidence model.
+When $\epsilon=0.35$ [Fig.
+15(a)], several agents are within each other’s confidence intervals, leading to active opinion exchange across multiple groups.
+This interaction results in partial convergence toward intermediate values after a few iterations.
+When $\epsilon=0.30$ [Fig.
+15(b)], the interaction range slightly decreases.
+However, since most of the key interaction links remain, the evolution results are the same as in the $\epsilon=0.35$ case.
+For $\epsilon=0.25$ [Fig.
+15(c)], the threshold is further reduced, causing only a subset of agents to interact.
+Opinion exchange occurs primarily within smaller local clusters, and the number of isolated agents increases.
+When $\epsilon=0.20$ [Fig.
+15(d)], $\epsilon=0.15$ [Fig.
+15(e)], and $\epsilon=0.10$ [Fig.
+15(f)], as the threshold gradually decreases, the information exchange between agents becomes less frequent, making it difficult for them to reach consensus in the end.
+Through the analysis, it is evident that this model is susceptible to the threshold, where even slight differences can result in significant variations in the results.
+V-C3 Heterogeneous HK Bounded Confidence Model
+The heterogeneous HK bounded confidence model extends the homogeneous version by assigning different threshold values to agents.
+In the first case, the threshold $\epsilon_{i}$ of each agent is listed below:
+[…]
+[…]
+[…]
+[…]
+[…]
+In the second case, $\epsilon_{10}$ is changed from 0.3 to 0.2, while all other values remain unchanged.
+In the third case, $\epsilon_{17}$ is changed from 0.7 to 0.3, while all other values remain unchanged.
+The results are shown in Fig.
+16.
+The final results are distinctly different.
+Case 1 converges to two distinct opinion clusters, Case 2 produces a different set of stable clusters, and Case 3 results in yet another opinion distribution.
+These results demonstrate that the model is also susceptible to individual parameter values, where even a small change can lead to significantly different results.
+Figure 14: Opinion Evolution under the DeGroot Model with Uniform and Distance-Based Weights.
+Figure 15: Opinion Evolution of All Agents under Different Confidence Bounds in the Homogeneous HK Model.
+Figure 16: Opinion Evolution of All Agents under Different Confidence Bounds in the Heterogeneous HK Model.
+V-D Potential Applications
+V-D1 Brain-Computer Interface (BCI) systems
 BCI systems often involve non-invasive brain signals such as electroencephalography (EEG) and functional near-infrared spectroscopy (fNIRS).
-These signals usually have high variability, strong individual differences, and incomplete information.
+These signals typically exhibit high variability, significant individual differences, and incomplete information.
 Traditional classification methods require making clear decisions for all states.
@@
 Figure 17: Description of Brain Regions and EEG-Based Paradigms. (a) Four Regions of the Cortex. (b) Four Other Common EEG-Based Paradigms [35].
-V-G2 Opinion Fusion in Smart City Governance
+V-D2 Opinion Fusion in Smart City Governance
 In smart city governance, such as public policy making, intelligent traffic management, and urban risk warning, a large amount of public opinion needs to be collected and integrated.
@@
 The proposed model allows individual opinions to be expressed in natural language and handles uncertain areas through deferred decisions.
-This makes the opinion fusion process more understandable and more tolerant of ambiguity.
-V-G3 Multi-Robot Coordination in Disaster Response and Agricultural Monitoring
-In scenarios such as post-disaster search and rescue and agricultural monitoring, multi-robot systems need to make collaborative decisions based on environmental sensing, their own abilities, and task requirements.
-Traditional assignment strategies often require each robot to immediately accept or reject a task.
+This makes the opinion fusion process more comprehensible and less prone to ambiguity.
+V-D3 Multi-Robot Coordination in Disaster Response and Agricultural Monitoring
+In scenarios such as post-disaster search and rescue and agricultural monitoring, multi-robot systems must make collaborative decisions based on environmental sensing, their own capabilities, and task requirements.
+Traditional assignment strategies often require each robot to accept or reject a task immediately.
 However, in uncertain environments or when task complexity is unclear, this approach may lead to poor task allocation or wasted resources.
 By introducing the 3WD theory, robots are allowed to delay task decisions, so they can make better choices when more information becomes available.
-At the same time, by adjusting the communication network dynamically, robots can improve their cooperation based on task similarity or path overlap, which increases the overall efficiency and robustness of the system.
+At the same time, by dynamically adjusting the communication network, robots can enhance their cooperation based on task similarity or path overlap, thereby increasing the overall efficiency and robustness of the system.
 VI Conclusion
@@
 By simulating the information exchange and dynamic negotiation among UAVs, we demonstrate that the mechanism is effective in practical distributed intelligent systems.
-The experimental comparisons further verify the advantages of the proposed algorithm.
+The experimental comparisons further verify the advantages of the algorithm.
 As AI agents powered by large language models (LLMs) play an increasingly important role in group decision-making, how to combine the diverse solutions generated by different models has become a key issue.
 Future research can focus on conflict resolution between models, trust evaluation mechanisms, and integration strategies to improve the overall level of collaborative intelligence in the system.
-Also, we will add reinforcement learning or adaptive weighting based on past interactions to enable agents to evolve their decision strategies over time.
+Additionally, we will incorporate reinforcement learning or adaptive weighting based on past interactions to enable agents to refine their decision strategies over time.